OpenStudy (anonymous):

Solve for x: 2 over 3x + 1 = 27

OpenStudy (anonymous):

2/3x+1=27 2+3x/3x=27 2+3x=27*3x 2+3x=81x 2=78x 1/39=x

OpenStudy (anonymous):

whole 3x+1 is over 2 or nly 3x?